# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ IV.2 Algorithm to Construct Boundary Stabilizers
where the $C(e)$ are edge terms required for the case of twisted quantum doubles (e.g. Eq.(23)). Here, the summation is over all vertices $V$, edges $E$, and plaquettes $P$ as the model is assumed to live on a closed surface $\Sigma$. And, as before, the star and plaquette operators are the standard tensor products of generalized Pauli operators; Eq. (6). The terms of this Hamiltonian generate a stabilizer group, which we denote by $\mathcal{S}_{H}$.
@@
We want to modify both the Hamiltonian and the stabilizer group to have a spatial boundary $\partial R$ bounding a region $R\subset\Sigma$ of the lattice. In addition, as boundaries are in 111-1 correspondence with Lagrangian subgroups (the anyons they condense), we have to choose the kind of boundary which corresponds to a Lagrangian subgroup $\mathcal{L}$. We do this in six steps:
First, we choose a $2D$ region $R\subset\Sigma$ in the lattice with a $1D$ boundary $\partial R$ and choose a Lagrangian subgroup $\mathcal{L}$ for the boundary. The type of the boundary is independent of its microscopic spatial geometry (one does not need to fix a specific shape of the boundary beforehand). The boundary can live on the direct or dual lattice, or a mixture of them. One choice for the boundary is shown by the pink line in Fig. 12.
@@
Sixth, restrict the stabilizers at the boundary to the uncondensed edges only. If, after this step, new qudits are condensed, then repeat step five until no more qudits are condensed  [^2] In this case, it is optional to remove these qudits or keep them. Keeping them will result in a boundary that strictly follows the boundary $\partial R$. We choose to erase all condensed qudits for clarity. The new Hamiltonian will be the summation of the generators of $\mathcal{S}_{H^{\prime}}$.
@@
Where $A,B$, and $C$ are the bulk stabilizers that do not overlap with $\partial R$. $J(v,p)$ represents a generic boundary term that can depend on a vertex and an adjacent plaquette at the same time. The algorithm is summarized in Alg.1.
@@
1:procedure Bdry($\mathcal{S}_{H}$)
2:  Choose a region $R$ with boundary $\partial R$ and a Lagrangian subgroup $\mathcal{L}_{R}$.
@@
7:   Erase condensed qudits.
8:   Restrict boundary stabilizers to their uncondensed support.
@@ IV.3 Examples
Example IV.1 (Smooth boundary for the $\mathbb{Z}_{4}$ surface code).
@@
This is a perfectly valid boundary for the surface code. The seemingly rough boundary is actually a smooth boundary because it condenses $m$ anyons. Had we chosen the pink path to lie in the dual lattice only, we would not have needed to deal with this, as exemplified by the other vertical pink line. Removing the edges that are now touched by only one $X$ operator gives us the (geometrically) smooth boundary [62] illustrated in Fig. 15(b). Sixth, we restrict the new stabilizers to their support on the uncondensed edges. Some of the vertices now have a star that is supported on only 3 edges.
@@
In example IV.1, the fourth step was trivial because of the simplicity of the surface code. In the next example, we treat the slightly more complicated case of a new boundary in $\mathbb{Z}_{4}$.
Example IV.2 (Even boundary for the $\mathbb{Z}_{4}$ surface code).
@@
Fifth, we now remove the condensed edges. Since a pair of degree two constraints is simultaneously applied to a four-dimensional degree of freedom, all edges in the bulk of the region $R$, or on the pink boundary $\partial R$, will condense to the vacuum. In general, the qudits inside $R$ or on $\partial R$ will always condense. We reach Fig. LABEL:fig:z4e2m24.
@@
Observe that the support of the new star $A^{2}(v)$ operators on the boundary $\partial R$ is reduced to a single edge. Additionally, the new plaquette operators $B^{2}(p)$ now have support on three edges. After restricting the stabilizers on the boundary to their uncondensed supports, no more qudits are condensed, and we are done. We arrive at Fig. LABEL:fig:z4e2m25 which describes the even boundary that condenses the $\{1,e^{2},m^{2},e^{2}m^{2}\}$ anyons. This gives the boundary in Eq. (14)
$H_{\mathbb{Z}_{4},\text{Even}}=H_{\mathbb{Z}_{4}}-\sum_{p,e\in\partial\Sigma}(B^{2}(p)+X^{2}(e))+H.C.$ (46)
@@
Example IV.3 (Boundary for the Doubled Semion (DS) phase).
@@
Fifth, we now remove the condensed edges. All edges in the bulk of the region $R$, or on the pink boundary $\partial R$, will condense to the vacuum. We reach Fig. LABEL:fig:dsbdry4.
@@
Observe that the support of the $X^{2}Z^{2}$ operators on the boundary $\partial R$ is reduced to a single edge. Additionally, the new star operators $A^{2}(v)$ now have support on three edges. After restricting the stabilizers on the boundary to their uncondensed supports, no more qudits are condensed, and we are done. We arrive at Fig. LABEL:fig:dsbdry5, which describes the unique DS boundary that condenses the $\{1,b\}$ anyons. This gives the boundary Hamiltonian:
$H_{DS-bdry}=H_{\text{DS}}-\sum_{v,e\in\partial\Sigma}(A^{2}(v)+X^{2}(e))+H.C.$ (49)
IV.4 Domain Walls
@@
Example IV.4 (Domain Wall $(e\leftrightarrow m)$ for $\mathbb{Z}_{4}$ (or $\mathbb{Z}_{n}$)).
@@
In the fifth step, we erase condensed qudits. These are all the qudits in the bulk of $R$ or on its two-component boundary $\partial_{1}R$ and $\partial_{2}R$. Note that the edges of the second layer that lie above $\partial_{1}R$ will not be condensed. After condensing qudits, we also removed all stabilizers that only act on condensed qudits for clarity, Fig. 29.
@@
The sixth step is to restrict the stabilizers at the boundary to their uncondensed support. This results in the following stabilizers:
@@
Figure 29: In the last step of the algorithm, we restrict the stabilizers to their uncondensed support. After unfolding, we have the twist operators $T(v,p)$ (purple) that live on a vertex and an adjacent plaquette, and also the $XZ(e)$ (ellipse) operators connecting two edges.
@@
Example IV.5 (Domain wall between $D(\mathbb{Z}_{4})$ and DS).
@@
Fifth, we now remove the condensed edges. In this case, no edges are condensed. Subsequently, the sixth step is also trivial. This example shows that in a certain situation, one does not need to do the folding. In App. A, we corroborate this by deriving the domain walls using the folding method.
V Calculating GSD using Stabilizers
To use Abelian (twisted) quantum doubles as quantum error-correcting codes, one must first compute the ground state degeneracy. This topological degeneracy corresponds to the dimension of the logical code space, and, to be used in fault-tolerant quantum computation, we must also understand how it is acted upon by logical operators. In this section, we detail how, within the stabilizer description, both local (e.g. stabilizer order) and global (e.g. boundary conditions) information is used to calculate the GSD.
The ground state projector has, by definition, the ground states in its image and all excited states in its kernel. Thus, the most straightforward method to compute the ground state degeneracy is to explicitly construct the ground state projector and take its trace [^3] The same logic applies for computing the dimensionality of any excited state sector.. However, due to the tensor product structure, this is not scalable in practice, and one must calculate the degeneracy using dimensional counting arguments defined by stabilizers, which, as detailed below, each simultaneously constrain the ground state degeneracy by their order.
Begin with an unconstrained Hilbert space of dimension $D$. We then consider the $n$ mutually commuting Hermitian stabilizer group generators $\left\{O_{i}\right\}$, each with order $m_{i}$ (i.e. $O^{m_{i}}_{i}=1$).
@@ V Calculating GSD using Stabilizers
Here, $n_{m}$ and $l_{m}$ are the numbers of generators and dependencies that have weight $m$, respectively. By definition, the $\operatorname{GSD}$ of topological theories does not depend on the system size nor its microscopic details, e.g., the particular triangulation of the surface. As we will see in the examples below, we can therefore use any minimal example to calculate the $\operatorname{GSD}$ and then extend the result for a larger or more complicated, but topologically equivalent, system.
V.1 Examples
Example V.1 ($D(\mathbb{Z}_{2})$ on a Disk I).
As a warm-up example, consider the $D(\mathbb{Z}_{2})$ code with the Hamiltonian $H=-\sum_{v\in V}A(v)-\sum_{p\in F}B(p)$. Here, $A(v)=\bigotimes_{j\in star(v)}X_{j}$ and $B(p)=\bigotimes_{m\in p}Z_{m}$ are order two operators $A(v)^{2}=B(p)^{2}=\mathbb{I}$. Consequently, we have the ground state projector $P_{\text{GS}}=(\prod_{v}\frac{1+A(v)}{2})(\prod_{p}\frac{1+B(p)}{2})$. The GSD computation is equivalent to taking the trace of the ground state projector as follows. Take the code to reside on a surface (i.e., open boundary conditions), as illustrated in Fig. 33, with one smooth and one rough boundary.
Figure 33: $D(\mathbb{Z}_{2})$ on a surface with rough and smooth boundaries meeting at only two points. Here, the number of rows and columns is three; $N=3$.
Assuming that we have $N$ rows and $N$ columns, and noting that there are two edges per unit cell, the total Hilbert space dimension is $D=2^{\#E}=2^{2N^{2}}$. To count the number of generating stabilizer constraints, we also define the number of plaquettes #P\#P and the number of vertices #V\#V. Since the plaquette and star operators are both order $2$, $n_{m=2}=\#P+\#V$. We have,
$\operatorname{GSD}$ $=\operatorname{Tr}(P_{\text{GS}})=\operatorname{Tr}\left(\prod_{v,p}(\frac{\mathbb{I}_{4}+A(v)}{2})(\frac{\mathbb{I}_{4}+B(p)}{2})\right)$ (56)
$=\operatorname{Tr}\left(\prod_{v,p}\frac{1}{2}{(\mathbb{I}_{4}+\bigotimes_{j\in s(v)}X_{j})}\frac{1}{2}(\mathbb{I}_{4}+\bigotimes_{m\in p}Z_{m})\right)$
$=\frac{1}{2^{\#V}2^{\#P}}\operatorname{Tr}\left(\prod_{v,p}{(\mathbb{I}_{4}+\bigotimes_{j\in s(v)}X_{j})}(\mathbb{I}_{4}+\bigotimes_{m\in p}Z_{m})\right)$
$=\frac{\operatorname{Tr}\left(\mathbb{I}_{2N^{2}}\right)}{2^{\#V}2^{\#P}}=\frac{2^{\#E}}{2^{\#V+\#P}}=\frac{2^{2N^{2}}}{2^{N^{2}}2^{N^{2}}}=\frac{D}{2^{2N^{2}}}=1$
In the second-to-last line, we used the property that tensor products of Pauli operators are traceless. In the last line, the identity on the total Hilbert space has trace $2^{\#E}$. The difference between this example (where $l_{2}=0$) and the typical toric code (where $l_{2}=2$) is that no product of stabilizers gives the global identity. This proves the ground state is non-degenerate and unique on an open disk with one smooth and one rough boundary.
If the code is re-gauged to add boundaries or domain walls as described in Alg. 1, we may sequentially update the GSD by only accounting for changes. Let us summarize the results of applying Alg. 1 as they relate to constraint dimensions and dependencies. This summary will culminate in equations to compute the updated $\operatorname{GSD}^{\prime}$ as a function of the old one. Suppose we know the $\operatorname{GSD}$ for a certain configuration realizing stabilizer group $\mathcal{S}_{H}$ with generators $s_{H}=\{O_{i}\}$ with $O_{i}^{m_{i}}=1$ and dependencies $l_{m}$ as from Eq. 56. We can apply the algorithm to obtain a different configuration with Hilbert space dimension $D^{\prime}$, stabilizer group $\mathcal{S}_{H^{\prime}}$, generators $s_{H^{\prime}}=\{O^{\prime}_{j}\}$, with $O^{m^{\prime}_{j}}_{j}=1$ and dependencies $l^{\prime}_{m}$. After the algorithm is complete, a set of new generators $O_{j}^{\prime}\in s_{H^{\prime}}$ will be added to the new Hamiltonian $H^{\prime}$:
@@ V.1 Examples
To get a feeling for Eq. 59, let us start with example V.1 and only change the boundary conditions to make the surface code.
Example V.2 ($\mathbb{Z}_{2}$ Surface Code).
Starting with the configuration of Example V.1 with $l_{2}=0$ and $\operatorname{GSD}=1$, we want to track changes that make the code into the surface code Fig. 34. Since all the operators here have order $2$, the change in $\operatorname{GSD}$ can be done by tracking the three numbers $r_{D}$, $l_{2}$, and $n_{2}$.
Figure 34: $D(\mathbb{Z}_{2})$ surface code. Compared with Example V.1, we see that the only difference, as far as the constraint counting is concerned, is the number of star operators, which decreased by $1$.
First, we have $r_{D}=1$. Second, all new operators are linearly independent, so $l^{\prime}_{2}=0$ and thus $\Delta l_{2}=l^{\prime}_{2}-l_{2}=0$. Note that the spatial placement of an operator does not matter. Even its exact action on the Hilbert space does not matter. For the purposes of the $\operatorname{GSD}$, the operator can only be labeled by its order and dependencies. Thus, tracking the change $\Delta n_{2}$ amounts to the difference of the sum of plaquettes and stars between the two examples. By direct counting, $\Delta n_{2}=-1$ since we lost one weight two operator (a star operator). We then use Eq. (59) to compute the new GSD:
$\operatorname{GSD}^{\prime}$ $=\frac{r_{D}}{\prod_{|m|=2}m^{\Delta n_{m}-\Delta l_{m}}}\operatorname{GSD}$ (60)
$=\frac{r_{D}}{2^{\Delta n_{2}-\Delta l_{2}}}\operatorname{GSD}=\frac{1}{2^{-1-0}}\times 1=2$
One can simply build on the surface code by introducing boundary conditions on the smooth edges, for example, to make a cylinder with two rough boundaries.
Example V.3 ($\mathbb{Z}_{2}$ on a cylinder with two rough boundaries I).

### v4

@@ V.3 Algorithm to Construct Boundary Stabilizers
where the $C(e)$ are edge terms required for the case of twisted quantum doubles (e.g. Eq.(20)). Here, the summation is over all vertices $V$, edges $E$, and plaquettes $P$ as the model is assumed to live on a closed surface $\mathcal{M}$. And, as before, the star and plaquette operators are the standard tensor products of generalized Pauli operators; Eq. (5). The terms of this Hamiltonian generate a stabilizer group, which we denote by $\mathcal{S}_{H}$.
@@
We want to modify both the Hamiltonian and the stabilizer group to have a spatial boundary $\partial R$ bounding a region $R\subset\mathcal{M}$ of the lattice. In addition, as boundaries are in 111-1 correspondence with Lagrangian subgroups (the anyons they condense), we have to choose the kind of boundary which corresponds to a Lagrangian subgroup $\mathcal{L}$. We do this in six steps:
First, we choose a $2D$ region $R\subset\mathcal{M}$ in the lattice with a $1D$ boundary $\partial R$ and choose a Lagrangian subgroup $\mathcal{L}$ for the boundary. The type of the boundary is independent of its microscopic spatial geometry (one does not need to fix a specific shape of the boundary beforehand). The boundary can live on the direct or dual lattice, or a mixture of them. One choice for the boundary is shown by the pink line in Fig. 16.
@@
Sixth, restrict the stabilizers at the boundary to the non-trivial qudits only. If, after this step, new qudits are condensed, then repeat step five until no more qudits are condensed  [^1] In this case, it is optional to remove these qudits or keep them.. Keeping them will result in a boundary that strictly follows the boundary $\partial R$. We choose to erase all trivialized qudits for clarity. The new Hamiltonian will be the summation of the generators of $\mathcal{S}_{H^{\prime}}$.
@@
Where $A,B$, and $C$ are the bulk stabilizers that do not overlap with $\partial R$. $J(v,p)$ represents a generic boundary term that can depend on a vertex and an adjacent plaquette at the same time. The algorithm is summarized in Alg.2.
@@
1:procedure Bdry($H,R,\mathcal{L}_{R}$)
2:A region $R$ with boundary $\partial R$ and a Lagrangian subgroup $\mathcal{L}_{R}$.
@@
7:   Erase trivialized qudits.
8:   Restrict boundary stabilizers to their non-trivialized support.
@@ V.4 Examples of Boundary Construction
Example V.1 (Smooth boundary for the $\mathbb{Z}_{4}$ surface code).
@@
This is a perfectly valid boundary for the surface code. The seemingly rough boundary is actually a smooth boundary because it condenses $m$ anyons. Had we chosen the pink path to lie in the dual lattice only, we would not have needed to deal with this, as exemplified by the other vertical pink line. Removing the edges that are now touched by only one $X$ operator gives us the (geometrically) smooth boundary [26] illustrated in Fig. 19(b). Sixth, we restrict the new stabilizers to their support on the non-trivial qudits. Some of the vertices now have a star that is supported on only 3 edges.
@@
In example V.1, the fourth step was trivial because of the simplicity of the surface code. In the next example, we treat the slightly more complicated case of a new boundary in $\mathbb{Z}_{4}$.
Example V.2 (Even boundary for the $\mathbb{Z}_{4}$ surface code).
@@
Fifth, we now remove the trivialized qudits. Since a pair of degree two constraints is simultaneously applied to a four-dimensional degree of freedom, all edges in the bulk of the region $R$, or on the pink boundary $\partial R$, will condense to the vacuum. In general, the qudits inside $R$ or on $\partial R$ will always condense. We reach Fig. LABEL:fig:z4e2m24.
@@
Observe that the support of the new star $A^{2}(v)$ operators on the boundary $\partial R$ is reduced to a single edge. Additionally, the new plaquette operators $B^{2}(p)$ now have support on three edges. After restricting the stabilizers on the boundary to their non-trivialized supports, no more qudits are condensed, and we are done. We arrive at Fig. LABEL:fig:z4e2m25 which describes the even boundary that condenses the $\{1,e^{2},m^{2},e^{2}m^{2}\}$ anyons. This gives the boundary in Eq. (14)
$H_{\mathbb{Z}_{4},\text{Even}}=H_{\mathbb{Z}_{4}}-\sum_{p,e\in\partial\mathcal{M}}(B^{2}(p)+X^{2}(e))+H.C.$ (41)
@@
Example V.3 (Boundary for the Doubled Semion (DS) phase).
@@
Fifth, we now remove the trivialized qudits. All edges in the bulk of the region $R$, or on the pink boundary $\partial R$, will condense to the vacuum. We reach Fig. LABEL:fig:dsbdry4.
@@
Observe that the support of the $X^{2}Z^{2}$ operators on the boundary $\partial R$ is reduced to a single edge. Additionally, the new star operators $A^{2}(v)$ now have support on three edges. After restricting the stabilizers on the boundary to their non-trivialized supports, no more qudits are condensed, and we are done. We arrive at Fig. LABEL:fig:dsbdry5, which describes the unique DS boundary that condenses the $\{1,b\}$ anyons. This gives the boundary Hamiltonian:
$H_{DS-bdry}=H_{\text{DS}}-\sum_{v,e\in\partial\mathcal{M}}(A^{2}(v)+X^{2}(e))+H.C.$ (44)
V.5 Examples of Domain Walls Construction
@@
Example V.4 (Domain Wall $(e\leftrightarrow m)$ for $\mathbb{Z}_{4}$ (or $\mathbb{Z}_{n}$)).
@@
In the fifth step, we erase trivialized qudits. These are all the qudits in the bulk of $R$ or on its two-component boundary $\partial_{1}R$ and $\partial_{2}R$. Note that the edges of the second layer that lie above $\partial_{1}R$ will not be condensed. After condensing qudits, we also removed all stabilizers that only act on trivialized qudits for clarity, Fig. 33.
@@
The sixth step is to restrict the stabilizers at the boundary to their non-trivialized support. This results in the following stabilizers:
@@
Figure 33: In the last step of the algorithm, we restrict the stabilizers to their non-trivialized support. After unfolding, we have the twist operators $T(v,p)$ (purple) that live on a vertex and an adjacent plaquette, and also the $XZ(e)$ (ellipse) operators connecting two edges.
@@
Example V.5 (Domain wall between $D(\mathbb{Z}_{4})$ and DS).
@@
Fifth, we now remove the trivialized qudits. In this case, no edges are condensed. Subsequently, the sixth step is also trivial. This example shows that in a certain situation, one does not need to do the folding. In App. A, we corroborate this by deriving the domain walls using the folding method.
VI Calculating Logical GSD
To confirm our Abelian (twisted) quantum doubles are quantum error-correcting codes, we follow Kitaev’s original logic 107 and rigorously compute the ground state degeneracy. This topological degeneracy corresponds to the dimension of the logical code space that is acted upon by logical operators as previously described in Sec. IV.
The ground state projector has, by definition, the ground states in its image and all excited states in its kernel. Thus, the most straightforward method to compute the ground state degeneracy is to explicitly construct the ground state projector and take its trace  [^2] The same logic applies for computing the dimensionality of any excited state sector. However, due to the tensor product structure, this is not scalable in practice.
In this section, we detail two complementary methods to efficiently calculate, modify, and reason about the $\operatorname{GSD}$ for composite topological codes. The first method utilizes the microscopic stabilizer description, where both local (e.g. stabilizer order) and global (e.g. boundary conditions) information is used to calculate the GSD. This is detailed below in Sec.VI. Alternatively, one can use a coarse-grained macroscopic description to compute the GSD using knowledge of logical ribbon operators and pants decomposition as detailed in Sec.VI.2, which follows Ref. 120.
VI.1 GSD using Microscopic Stabilizers
In this subsection, we assume we are given a concrete stabilizer layout for a system, and then compute its GSD. Begin with an unconstrained Hilbert space of dimension $D$. In the simple and most common case, all qudits will have the same dimension $d$ and thus, $D=d^{\#E}$ where #E\#E is the number of edges in the system, each supporting a single qudit [1]. We then consider the $n$ mutually commuting stabilizer group generators $\left\{O_{i}\right\}$, each with order $m_{i}$ (i.e. $O^{m_{i}}_{i}=1$). Note that, in general, for higher-dimensional qudits ($d>2$), these generators do not have to be Hermitian.
@@ VI.1 GSD using Microscopic Stabilizers
Here, $n_{m}$ and $l_{m}$ are the numbers of generators and dependencies that have order $m$, respectively.
By definition, the $\operatorname{GSD}$ of topological theories does not depend on the system size nor its microscopic details, e.g., the particular triangulation of the surface. As we will see in the examples below, we can therefore use any minimal example to calculate the $\operatorname{GSD}$ and then extend the result for a larger or more complicated, but topologically equivalent, system.
Updating GSD–If the code is re-gauged to add boundaries or domain walls, as described in Alg. 2, we may sequentially update the GSD by only accounting for changes. Let us summarize the results of applying Alg. 2 as they relate to constraint dimensions and dependencies. This summary will culminate in equations to compute the updated $\operatorname{GSD}^{\prime}$ as a function of the old one. Suppose we know the $\operatorname{GSD}$ for a certain configuration realizing stabilizer group $\mathcal{S}_{H}$ with generators $s_{H}=\{O_{i}\}$ with $O_{i}^{m_{i}}=1$ and dependencies $l_{m}$. We can apply the algorithm to obtain a different configuration with Hilbert space dimension $D^{\prime}$, stabilizer group $\mathcal{S}_{H^{\prime}}$, generators $s_{H^{\prime}}=\{O^{\prime}_{j}\}$, with $O^{m^{\prime}_{j}}_{j}=1$ and dependencies $l^{\prime}_{m}$. After the algorithm is complete, a set of new generators $O_{j}^{\prime}\in s_{H^{\prime}}$ will be added to the new Hamiltonian $H^{\prime}$:
@@
VI.1.1 Examples using Minimal Stabilizer Models
To get a feeling for Eq. 53, let us start with example VI.1 where we modify the surface code in Fig. 37 with $\operatorname{GSD}=2$ to make a cylinder with two rough boundaries.
Figure 37: $D(\mathbb{Z}_{2})$ surface code with opposite smooth and rough boundary conditions, and which has $\operatorname{GSD}=2$.
Example VI.1 ($\mathbb{Z}_{2}$ on a cylinder with two rough boundaries).